# LLINBO: Trustworthy LLM-in-the-Loop Bayesian Optimization

### v1

@@ 2.4 LLINBO-Justify: Surrogate-driven Rejection of LLM’s Suggestions
$\alpha_{\text{UCB}}(x_{\text{LLM},t},F_{t-1})\leq\max_{x}\alpha_{\text{UCB}}(x,F_{t-1})-\psi_{t},$
@@
An upper bound on the cumulative regret for LLINBO-Justify is provided in Theorem 2. We observe that, regardless of whether $x_{\text{LLM},t}$ is accepted or not, the next query design $x_{t}$ (either $x_{\text{LLM},t}$ or $x_{\mathcal{GP},t}$) always lies within the $\psi_{t}$-suboptimal region of $\alpha_{\text{UCB}}(x,F_{t-1})$. Leveraging this observation along with classical UCB analysis techniques [32], the result follows directly.
Theorem 2 (Proof in Appendix A.2).
Suppose that Assumptions 1-2 hold, $\psi_{t}\in\mathcal{O}(1/\sqrt{t})$, and LLINBO-Justify is applied to generate $x_{t}\ \forall t\in[T]$, then the cumulative regret is upper bounded by
$R_{T}=\sum_{t=1}^{T}r_{t}\leq\sum_{i=1}^{T}\delta_{t}+2\beta_{T}\sum_{i=1}^{T}\sigma_{t-1}(x_{t})=\mathcal{O}(\sqrt{T})+\beta_{T}\mathcal{O}(\sqrt{T\gamma_{T}}).$
@@ 2.5 LLINBO-Constrained: Constrain Surrogates on LLM’s Suggestions
This essentially leads to a constrained $\mathcal{GP}$, a $\mathcal{CGP}$. While $\mathcal{CGP}$ does not admit a closed-form posterior, one can readily draw function realizations from it via rejection sampling and approximate the AF using Monte Carlo (MC) [6].
@@
Fig. 2 illustrates the behavior of LLINBO-Constrained. Critically, more output samples are retained when the constraint is satisfied, reflecting posterior support for $x_{\text{LLM},t}$ as a high-quality candidate. In such cases, the mean function under the updated surrogate $F_{t-1}^{+}$ becomes elevated near $x_{\text{LLM},t}$, highlighting promising regions for subsequent exploration (see Fig. 2(a)–(b)). Conversely, when $x_{\text{LLM},t}$ strongly contradicts the current posterior, no samples are retained ($|I_{t}|=0$), and the surrogate remains unchanged, i.e., $F_{t-1}=F_{t-1}^{+}$, effectively discarding $x_{\text{LLM},t}$ in favor of $x_{\mathcal{GP},t}$ extracted by solving (1) (see Fig. 2(c)–(d)). This selective retention mechanism is key to maintaining the trustworthiness of the BO process and underpins the theoretical guarantees discussed later.
@@
(b) Two $\mathcal{GP}$s (blue and green curves and shaded areas) are constructed based on the union of each retained sample and $\mathcal{D}_{t-1}$.
(c) All points lie below than $\kappa_{t-1}$ (no retained points).
@@
Theorem 3 (Proof in Appendix A.3).
@@
where $\tilde{\beta}_{t}$ is given by
$\tilde{\beta}_{t}=2B+2R\sqrt{2\left(\gamma_{t}+1+\ln\left(\tfrac{4T}{\delta}\right)\right)}+\sqrt{2\ln\left(\tfrac{4S_{t}T}{\delta}\right)}.$
Compared to Assumption 2, Theorem 3 includes an additional term involving $S_{t}$, reflecting the cost of sampling uncertainty. As $S_{t}$ grows, the potential for deviation increases, requiring a larger $\tilde{\beta}_{t}$ to maintain the same confidence level. As such, Theorem 3 builds a uniform high-probability bound between the posterior mean of the $\mathcal{CGP}$ and $f$. With this, Theorem 4 then upper bounds the cumulative regret for LLINBO-Constrained.
Theorem 4 (Proof in Appendix A.3).
Assume the conditions for Theorem 3 hold and suppose $S_{t}\in\mathcal{O}(1/t)$, and LLINBO-Constrained is used to generate $x_{t}$ for all $t\in[T]$. Then, $R_{T}$ satisfies
@@ 3 Numerical Studies
We evaluate the proposed methods on two core BO tasks: black-box optimization and hyperparameter tuning, using two representative benchmarks: BO and LLAMBO, the most recent state-of-the-art framework introduced by Liu et al. [25]. While effective, implementing LLAMBO can be computationally expensive due to the extensive prompting required to generate multiple candidate designs and surrogate evaluations. To mitigate this overhead, we develop LLAMBO-light, a lightweight alternative that directly prompts the LLM with the problem context and historical observations to produce the next evaluation design. LLAMBO-light serves both as the embedded LLM agent within our proposed three mechanisms and as a baseline. We should note that this is still an emerging area with limited prior work. For a full description of our experimental setup and prompt designs, see Appendix B.
@@
For each task with a $D$-dimensional design space, we generate an initial dataset $\mathcal{D}_{0}$ with $D$ observations. This is done via prompting within the problem context, also known as warmstarting, for methods that utilize LLMs (ours, LLAMBO, LLAMBO-light), and via random sampling for BO. To capture the uncertainty in each method’s performance, we perform a total of 10 replications. We use UCB as the AF, and set the relevant parameters as follows: $p_{t}=\min(t^{2}/T,1)$, $S_{t}=10^{4}/t^{2}$, $\psi_{t}=\frac{1}{t}\sigma_{0}(x_{\text{LLM},1})$ and $\beta_{t}=2\text{log}\frac{tD\pi^{2}}{0.1*6}$ (as shown effective by Srinivas et al. [32]).
Black-box optimization.
@@
[…]
Figure 3: $G_{t}$ comparison for black-box optimization. Each line shows the mean regret, shaded with 95% confidence intervals. Proposed methods:     LLINBO-Transient,     LLINBO-Justify,     LLINBO-Constrained. Baselines:   LLAMBO,   LLAMBO-light,   BO.
Fig. 3 shows the regret curves for all methods across the six benchmark functions. Based on these results, we highlight several key insights. First, and perhaps most evidently, LLM-assisted BO (LLAMBO-light and LLAMBO) significantly underperform compared to other benchmarks. In many cases, their regret curves remain flat, especially in higher dimensions. This supports our motivation: LLMs can assist with black-box optimization but are not yet reliable as standalone agents. Second, methods involving LLMs, including ours, achieve a strong early lead. This suggests that LLMs can effectively leverage problem context to quickly identify promising regions, making them a useful complement to BO frameworks. Third, we observe that our hybrid mechanisms consistently outperform the benchmarks across all functions. This superiority is especially evident in the early rounds and gradually diminishes as more data is collected. This trend is not surprising; statistical surrogate models become more accurate with additional data, aligning with our core philosophy of reducing reliance on LLMs as the optimization process progresses.
Hyperparameter tuning.
[…]
Figure 4: MSE comparison for hyperparameter tuning. Each line shows the mean MSE, shaded with 95% confidence intervals. Proposed methods:     LLINBO-Transient,     LLINBO-Justify,     LLINBO-Constrained. Baselines:   LLAMBO,   LLAMBO-light,   BO.
4 Application to 3D Printing
In addition to the numerical evaluation above, we further assess the performance of our method through a case study in 3D printing, aimed at reducing stringing in a printed product. Stringing (Fig. 5(b)) is a prevalent defect in fused filament fabrication (FFF) 3D printing. FFF is commonly used for rapid prototyping and low-cost part production. However, stringing degrades surface quality and often requires additional post-processing [26]. This study aims to optimize the design parameters of a Creality Ender 3 desktop FFF printer (Fig. 5(a)), including nozzle temperature, Z hop height, retraction distance, outer wall wipe distance, and coasting volume, using stringing percentage as the outcome variable. Further details about the parameters can be found in Appendix C.
Experiment setup. All experiments were conducted on a single printer using PETG filament [16], selected for its high tendency to produce stringing (see Fig.5(b)). We adopted a standard two-column geometry with a horizontal gap, commonly used in stringing evaluations [15]. At each iteration, after printing the object with the proposed parameters, the stringing percentage (ranging from 0 to 100%) was quantified by converting the part’s image to grayscale and calculating the ratio of bright pixels within a predefined region between the columns (details in Appendix C.1).
Due to the cost associated with this experiment (each run takes several hours), we limit our comparison to LLINBO-Transient with $p_{t}=1-\frac{1}{t}$, evaluated against LLAMBO, LLAMBO-light, and BO. All other settings follow Sec. 3. The prompts specifying the problem context and controllable parameters are provided in Appendix C.2. Unlike Sec. 3, the objective here is not exhaustive evaluation, but to demonstrate the effectiveness of our method and the broader potential of LLMs in optimal design.
The results are presented in Fig. 5(c), from which we draw several insights: (i) Our approach, LLINBO-Transient, demonstrates strong overall performance and ultimately achieves near-zero stringing. (ii) Methods utilizing LLMs achieve a strong head start compared to BO, highlighting the value of LLMs in optimal design, where they act as domain-aware agents that leverage contextual knowledge to warmstart the optimization process. (iii) Consistent with our simulation results, LLAMBO and LLAMBO-light perform poorly and do not exhibit a decreasing trend in regret, underscoring the risks of LLM-assisted BO. (iv) While BO shows improvement over time, our hybrid approach outperforms it. This again highlights the collaboration benefits between LLMs and surrogate experts.
5 Conclusion
To the best of our knowledge, this is the first framework for collaborative black-box optimization that integrates LLMs with statistical surrogates such as $\mathcal{GP}s$. It leverages LLMs’ contextual reasoning to generate high-quality designs early, while surrogate models refine and guide the search as data accumulates. The mechanisms developed under this framework exhibit strong performance, as demonstrated by both simulation and real-world case studies. While the use of LLMs in optimization remains in its infancy, we believe this line of research holds great promise for enabling more adaptive, data-efficient, and practical optimization strategies across a wide range of applications. The strength of our hybrid framework depends on parameters that are sensitive to how well the LLM understands the problem context in early stages. A promising direction is to link these parameters to a metric that quantifies an LLM understanding. Our overarching framework can potentially help design LLM-assisted optimization beyond black-box settings.
Appendix A Technical Results
@@
(Theorem 2 of [8]) Under Assumption 1 and 2, and let $\hat{\lambda}_{t}=1+2/t$. For arbitrary $\delta\in(0,1)$, with probability at least $1-\delta$, we have:
@@
Lemma 2 (Theorem 3 in [8]).
@@
Lemma 3 (Lemma 4 in Appendix of [8]).
@@
A.1 Proof of Theorem 1
The proof builds on the approach of [9], which uses the Azuma-Hoeffding inequality to derive a high-probability upper bound on the regret, transforming the expected regret into a probabilistic guarantee. Recall that when LLINBO-Transient is applied, $x_{t}$ is selected as

### v2

@@ 2.4 LLINBO-Justify: Surrogate-driven Rejection of LLM’s Suggestions
$\alpha_{\text{UCB}}(x_{\text{LLM},t},F_{t-1})\leq\max_{x}\alpha_{\text{UCB}}(x,F_{t-1})-\psi_{t},$ (2)
@@
An upper bound on $R_{T}$ for LLINBO-Justify is provided in Theorem 2. From (2), we observe that, regardless of whether $x_{\text{LLM},t}$ is accepted or not, the next query design $x_{t}$ (either $x_{\text{LLM},t}$ or $x_{\mathcal{GP},t}$) always lies within the $\psi_{t}$-suboptimal region of $\alpha_{\text{UCB}}(x,F_{t-1})$. Leveraging this observation along with classical UCB analysis techniques in Srinivas et al. (2009), the result follows directly.
Theorem 2 (Proof in Appendix B.2).
Suppose that Assumptions 1-2 hold and $\psi_{t}\in\mathcal{O}(1/\sqrt{t})$. Then, with probability at least $1-\delta$, $R_{T}$ is upper bounded by
$R_{T}=\sum_{t=1}^{T}r_{t}\leq\sum_{i=1}^{T}\psi_{t}+2\beta_{T}\sum_{i=1}^{T}\sigma_{t-1}(x_{t})=\mathcal{O}(\sqrt{T})+\beta_{T}\mathcal{O}(\sqrt{T\gamma_{T}}).$
@@ 2.5 LLINBO-Constrained: Constrain Surrogates on LLM’s Suggestions
This essentially leads to a constrained $\mathcal{GP}$, a $\mathcal{CGP}$. While $\mathcal{CGP}$ does not admit a closed-form posterior, one can readily draw function realizations from it via rejection sampling and approximate the AF using Monte Carlo (MC, Chen et al. (2025)).
@@
Fig. 2 illustrates the behavior of LLINBO-Constrained. Critically, more output samples are retained when the constraint is satisfied, reflecting posterior support for $x_{\text{LLM},t}$ as a high-quality candidate. In such cases, the mean function under the updated surrogate $F_{t-1}^{+}$ becomes elevated near $x_{\text{LLM},t}$, highlighting promising regions for subsequent exploration (see Fig. 2(a)–(b)). Conversely, when $x_{\text{LLM},t}$ strongly contradicts the current posterior, no samples are retained ($|I_{t}|=0$), and the surrogate remains unchanged, i.e., $F_{t-1}=F_{t-1}^{+}$, effectively discarding $x_{\text{LLM},t}$ in favor of $x_{\mathcal{GP},t}$ (see Fig. 2(c)–(d)). This selective retention mechanism is key to maintaining the trustworthiness of the BO process and underpins the theoretical guarantees discussed later.
@@
(b) Two $\mathcal{GP}$s (red and green curves and shaded areas) are constructed based on the union of each retained sample and $\mathcal{D}_{t-1}$.
(c) All points lie below than $\kappa_{t-1}$.
@@
Theorem 3 (Proof in Appendix B.3).
@@
where $\tilde{\beta}_{t}$ is given by $\tilde{\beta}_{t}=2B+2R\sqrt{2\left(\gamma_{t}+1+\ln\left(\tfrac{4T}{\delta}\right)\right)}+\sqrt{2\ln\left(\tfrac{4S_{t}T}{\delta}\right)}.$
Compared to Assumption 2, Theorem 3 includes an additional term involving $S_{t}$, reflecting the cost of sampling uncertainty. As $S_{t}$ grows, the potential for deviation increases, requiring a larger $\tilde{\beta}_{t}$ to maintain the same confidence level. As such, Theorem 3 builds a uniform high-probability bound between the posterior mean of the $\mathcal{CGP}$ and $f$. With this, Theorem 4 then upper bounds $R_{T}$ for LLINBO-Constrained.
Theorem 4 (Proof in Appendix B.3).
Assume the conditions for Theorem 3 hold and suppose $S_{t}\in\mathcal{O}(1/t)$. Then, with probability at least $1-\delta$, $R_{T}$ satisfies
@@ 3 Numerical Studies
We evaluate the proposed methods on two core BO tasks: BBO and HPT, using two representative benchmarks: BO and LLAMBO, the most recent state-of-the-art framework introduced by Liu et al. (2024). While effective, implementing LLAMBO can be computationally expensive due to the extensive prompting required to generate multiple candidate designs and surrogate evaluations. To mitigate this overhead, we develop LLAMBO-light, a lightweight alternative that directly prompts the LLM with the problem context and historical observations to produce the next evaluation design. LLAMBO-light serves both as the embedded LLM agent within our proposed three mechanisms and as a baseline. We should note that this is still an emerging area with limited prior work. For a full description of our experimental setup and prompt designs, see Appendix E.
@@
For each task with a $D$-dimensional design space, we generate an initial dataset $\mathcal{D}_{0}$ with $D$ observations. This is done via prompting within the problem context, also known as warmstarting, for methods that utilize LLMs (ours, LLAMBO, LLAMBO-light), and via random sampling for BO. To capture the uncertainty in each method’s performance, we perform a total of 10 replications. We use UCB as the AF, and set the relevant parameters as follows: $p_{t}=\min(t^{2}/T,1)$, $S_{t}=10^{4}/t^{2}$, $\psi_{t}=\frac{1}{t}\sigma_{0}(x_{\text{LLM},1})$ and $\beta_{t}=2\text{log}\frac{tD\pi^{2}}{0.1*6}$ (as shown effective by Srinivas et al. (2009)).
BBO task.
We utilize six commonly used simulation functions: Levy-2DD, Rastrigin-2DD, Branin-2DD, Bukin-2DD, Hartmann-4DD, and Ackley-6DD from Surjanovic & Bingham (2013). For each function, its characteristic patterns and the objective of the problem are incorporated into the prompts as part of the problem context (see Appendix E.1). Performance is reported in terms of the best observed regret, defined as $G_{t}=f(x^{*})-y^{*}_{t}$, where $y^{*}_{t}$ is the best outcome observed up to time $t$, and $f(x^{*})$ denotes the true global maximum. The total budget is set to $T=10D$.
(a) Branin-2DD
(b) Ackley-6DD
(c) Hartmann-4DD
(d) Levy-2DD
(e) Bukin-2DD
(f) Rastrigin-2DD
Figure 3: $G_{t}$ comparison for BBO. Each line shows the mean regret, shaded with 95% confidence intervals. Proposed methods:     LLINBO-Transient,     LLINBO-Justify,     LLINBO-Constrained. Baselines:   LLAMBO,   LLAMBO-light,   BO.
Fig. 3 shows the regret curves for all methods across the six benchmark functions. Based on these results, we highlight several key insights. First, and perhaps most evidently, LLM-assisted BO (LLAMBO-light and LLAMBO) significantly underperform compared to other benchmarks. In many cases, their regret curves remain flat, especially in higher dimensions. This supports our motivation: LLMs can assist with BBO but are not yet reliable as standalone agents. Second, methods involving LLMs, including ours, achieve a strong early lead. This suggests that LLMs can effectively leverage problem context to quickly identify promising regions, making them a useful complement to BO frameworks. Third, we observe that our hybrid mechanisms consistently outperform the benchmarks across all functions. This superiority is especially evident in the early rounds and gradually diminishes as more data is collected. This trend is not surprising; statistical surrogate models become more accurate with additional data, aligning with our core philosophy of reducing reliance on LLMs as the optimization process progresses.
HPT task.
We consider two physical simulation functions: the piston (Kenett & Zacks (1998)) and robot arm (An & Owen (2001)), along with three regression models: Random Forest (RF-4DD), Support Vector Regression (SVR-3DD), and XGBoost (XGB-4DD). The total budget is set to $T=5D.$ For each simulation function, we generate 1,000 data points and define the regret as the best-observed Mean Squared Error (MSE) at each iteration, where the MSE is obtained by fitting the corresponding regression model and evaluating it via 10-fold cross-validation. A detailed description of each data–regression model pair, along with the corresponding problem formulation, is provided in the prompt (see Appendix E.2). The results in Fig. 4 once again confirm the insights from the BBO task. Namely, we find that LLMs are often capable of generating high-quality designs in the early iterations by leveraging the problem context. Furthermore, our proposed LLM-$\mathcal{GP}$ collaborative mechanisms yield significantly lower MSE compared to all benchmarks across the tasks.
(a) Piston with RF-4DD
(b) Piston with SVR-3DD
(c) Piston with XGB-4DD
(d) Robot with RF-4DD
(e) Robot with SVR-3DD
(f) Robot with XGB-4DD
Figure 4: MSE comparison for HPT. Each line shows the mean MSE, shaded with 95% confidence intervals. Proposed methods:     LLINBO-Transient,     LLINBO-Justify,     LLINBO-Constrained. Baselines:   LLAMBO,   LLAMBO-light,   BO.
@@
While all proposed methods perform well in both BBO and HPT tasks, the choice among them ultimately depends on practical needs. We provide details for choosing among them and selecting hyperparameters in Appendix D.
4 Application to 3D Printing
In addition to the numerical evaluation, we further assess the performance of our method through a case study in 3D printing, aimed at reducing stringing in a printed product. Stringing (Fig. 5(b)) is a prevalent defect in fused filament fabrication (FFF) 3D printing. FFF is commonly used for rapid prototyping and low-cost part production. However, stringing degrades surface quality and often requires additional post-processing (Paraskevoudis et al. (2020)). This study aims to optimize the design parameters of a Creality Ender 3 desktop FFF printer (Fig. 5(a)), including nozzle temperature, Z hop height, retraction distance, outer wall wipe distance, and coasting volume, using stringing percentage as the outcome variable. Further details about the parameters can be found in Appendix F.
Experiment setup. All experiments were conducted on a single printer using PETG filament (Holcomb et al. (2022)), selected for its high tendency to produce stringing (see Fig.5(b)). We adopted a standard two-column geometry with a horizontal gap, commonly used in stringing evaluations (Haque (2020)). At each iteration, after printing the object with the proposed parameters, the stringing percentage (ranging from 0 to 100%) was quantified (details in Appendix F.1).
Due to the cost associated with this experiment (each run takes several hours), we limit our comparison to LLINBO-Transient with $p_{t}=1-\frac{1}{t}$, evaluated against LLAMBO, LLAMBO-light, and BO. All other settings follow Sec. 3. The prompts specifying the problem context and controllable parameters are provided in Appendix F.2. Unlike Sec. 3, the objective here is not full evaluation, but to demonstrate the effectiveness of our method and the broader potential of LLMs in optimal design.
Several insights can be draw from the results shown in Fig. 5(c): (i) Our approach demonstrates strong overall performance and ultimately achieves near-zero stringing. (ii) Methods utilizing LLMs achieve a strong head start compared to BO, highlighting the value of LLMs in optimal design. (iii) Consistent with our simulation results, LLAMBO and LLAMBO-light perform poorly and do not exhibit a decreasing trend in regret. (iv) While BO shows improvement over time, our hybrid approach outperforms it. This again highlights the collaboration benefits between LLMs and surrogate experts.
5 Conclusion
The proposed LLINBO framework leverages LLMs’ contextual reasoning to generate high-quality designs early, while surrogate models refine and guide the search as data accumulates. The mechanisms developed under this framework exhibit strong performance, as demonstrated by both simulation and real-world case studies. While the use of LLMs in optimization remains in its infancy, we believe this line of research holds great promise for enabling more adaptive, data-efficient, and practical optimization strategies across a wide range of applications. The strength of our hybrid framework depends on parameters that are sensitive to how well the LLM understands the problem context in early stages. A promising direction is to link these parameters to a metric that quantifies an LLM understanding. Our overarching framework can potentially help design LLM-assisted optimization beyond black-box settings.
Appendix A More Related Works
LLM-assisted BO.
Recently, with the few-shot learning ability of LLMs to generate high-quality answers from limited input, leveraging LLMs in the BO process has emerged as a promising yet relatively new research direction. For example, Liu et al. (2025) employed LLMs to solve multi-objective optimization problems, while Guo et al. (2024) extended their use to a broader set of tasks, including combinatorial optimization. More recently, Song et al. (2024) explored how LLMs can enhance BBO by leveraging textual knowledge and sequence modeling to improve generalization. These works highlight the potential of LLMs in various optimization settings, a direction that remains actively under investigation.
F-BO
To enable collaboration between LLMs and statistical surrogates in enhancing BO, while leveraging the distinctive ability of LLMs to provide a set of designs, we draw inspiration from the literature on F-BO. Federated learning (FL) aims to establish a collaborative framework that allows clients to work together while keeping their own data private. This setting has directly influenced prior work in F-BO, where a single design is often shared across clients.
For example, Yue et al. (2025) developed a consensus framework for collaborative BO, where the next design to query is selected as a weighted combination, dictated by a dynamically coupled stochastic consensus matrix, of the AF maximizers from all clients in the system, including each client’s own. Other works like Chen et al. (2025) and Dai et al. (2020) require only the design from other clients, while the former requires a design point from other clients directly, and the latter requires Random Fourier Features (Rahimi & Recht (2007)). A recent review on federated and collaborative BO can be found in Al Kontar (2024).
HAIC-BO
@@ Appendix A More Related Works
Appendix B Technical Results
@@
(Theorem 2 of Chowdhury & Gopalan (2017)) Under Assumption 1 and 2, and let $\hat{\lambda}_{t}=1+2/t$. For arbitrary $\delta\in(0,1)$, with probability at least $1-\delta$, we have:
@@
Lemma 2 (Theorem 3 in Chowdhury & Gopalan (2017)).
@@
Lemma 3 (Lemma 4 in Appendix of Chowdhury & Gopalan (2017)).
@@
B.1 Proof of Theorem 1
The proof builds on the approach of Dai et al. (2020), which uses the Azuma-Hoeffding inequality to derive a high-probability upper bound on the regret, transforming the expected regret into a probabilistic guarantee. Recall that when LLINBO-Transient is applied, $x_{t}$ is selected as